Calculate the pH of a Buffer solution formed by mixing 65.00 mL of 0.200 M NaHCO3 with 75 mL of 0.150 M Na2CO3
Answer #1

The chemical reaction can be given as

CO32- (aq) + H2O (l) ↔ HCO3- (aq) + OH- (aq)

Volume of NaHCO3 = 65.00 mL = 0.065 L

Molarity of NaHCO3 = 0.200 M

Volume of Na2CO3 = 75.00 mL = 0.075 L

Molarity of Na2CO3 = 0.150 M

Total volume = 0.065 L +0.075 L = 0.14 L

Concentration of NaHCO3 = (0.200) (0.065 L) /0.14L = 0.092 M

Concentration of Na2CO3 = (0.150) (0.075 L) /0.14L = 0.080 M

pH = pKa + log[CO32-]/ [HCO3- ]

= - log 15.6 x 10-11 + log (0.080/0.092)

= - log 5.6 x 10-11 + log 0.869

= - (-10.25) + (-0.06) = 10.25-.06 = 10.19